OpenStudy (anonymous):

2^(x+1)=3^x Help please. I got it down to (X+1)ln2=xln3 But now don't know what to do

OpenStudy (anonymous):

solve, treating \(\ln(2)\) and \(\ln(3)\) as constants

OpenStudy (anonymous):

\[(x+1)\ln(2)=x\ln(3)\] multiply out on the left get \[\ln(2)x+\ln(2)=\ln(3)x\] put all term with \(x\) on the right get \[\ln(2)=\ln(3)x-\ln(2)x\] factor out an \(x\) get \[\ln(2)=(\ln(3)-\ln(2))x\] and finally divide to get \[\frac{\ln(2)}{\ln(3)-\ln(2)}=x\]

OpenStudy (anonymous):

Wow. Thank you. I did not realize to distribute the ln2

OpenStudy (anonymous):

yw